OpenStudy (anonymous):

Find three solutions of 9x – 2y = 18.

sam (.sam.):

you're asking Find three solutions of 9x – 2y = 18.? 1....(0, 0), (9, 2), (2, 9) 2....(0, -2), (9, 0), (9, -2) 3....(0, 2), (-9, 0), (1, 7) 4.....(0, -9), (2, 0), (1, -4.5)?

sam (.sam.):

inserting the values (x,0) in the equation you get 9x-0=18 x=18/9=2 so (2,0) will be one solution inserting the values (0,y) in the equation you get 0-2y=18 y= -18/2= -9 so (0,-9) will be another solution now assume x=1 and find the value of y 9*1-2y=18 -2y=18-9 -2y=9 y= -9/2 y= -4.5 so (0,- 4.5) would be one solution answer 4.....(0, -9), (2, 0), (1, -4.5) is the correct option

OpenStudy (anonymous):

thank youu!!